OpenStudy (anonymous):

Multiply: [c+1/4][c+1/2] = Simplify answer??????

OpenStudy (anonymous):

I got c^2+3/4c+1/8????

OpenStudy (anonymous):

\[(c+\frac{1}{4})(x+\frac{1}{2})=c^2+\frac{1}{2}c+\frac{1}{4}c+\frac{1}{8}\] \[=c^2+\frac{3}{4}c+\frac{1}{8}\]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

like heck you do!

OpenStudy (anonymous):

if you multiply by 8 you get something 8 times as large!

OpenStudy (anonymous):

\[\frac{3}{4}\neq\frac{3}{4}\times 8\]

OpenStudy (anonymous):

I think the way you wrote it at first in fraction form is the way they want it

OpenStudy (anonymous):

if you are solving an equation you can multiply both sides by 8. or whatever. but if you just have an expression you cannot multiply by 8!

OpenStudy (anonymous):

it is not a matter of 'fraction form' take the expression \[x+3\] multiply it by 8 you get \[8x+24\] which is not the same at all.

OpenStudy (anonymous):

Oh ok gotcha, I see what you are saying

OpenStudy (anonymous):

whew. so don't do it! unless of course you are solving an equation and multiply both sides by something.